A manufacturer knows that their items have a normally distributed lifespan, with a mean of 14.9 years, and standard deviation of 2.1 years.

If you randomly purchase one item, what is the probability it will last longer than 19 years?
Answer #1

Solution :

Given that ,

mean = = 14.9

standard deviation = = 2.1

P(x > 19) = 1 - P(x < 19)

= 1 - P[(x - ) / < (19 - 14.9) / 2.1)

= 1 - P(z < 1.9524)

= 1 - 0.9746

= 0.0254

Probability = 0.0254